# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v3

@@ 2 Multi-orbital TPSC
$=\Big{(}D_{o_{4^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$ (8)
$\quad+C_{o_{1},o_{3^{\prime}}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$
$\quad+P_{o_{2^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}\Big{)}\,.$
@@
$\Big{(}D_{o_{4^{\prime}},o_{1}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})}$
$\quad-C_{o_{1},o_{3^{\prime}}}\sum_{s_{4^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})}$
@@
By combining Eq. (9) and Eq. (10) we obtain the explicit form of the Ansätze. It should be noted that there is an ambiguity in the way we define the prefactors: we can either define them as objects depending on one or on two spin indices. In the former case, we end up with three independent vertices, while in the latter case, we end up with five. The ansatz proportional to $P$ always depends only on a single spin. The origin of this ambiguity is the freedom of pulling the sum over the internal spin in Eq. (8) into the Ansatz (leading to TPSC3) or performing the Ansatz inside the sum (leading to TPSC5). Whether both of these choices are valid in the sense that no artificial $SU(2)$ symmetry breaking is introduced has to be checked by deriving the interaction vertex induced by them. Again we stress that there is technically a much larger group of possible Ansätze due to the freedom of shifting contributions to different interaction terms.
@@
In these equations, we introduced a short hand notation which allows us to drop redundant arguments, i.e. $n_{o_{1},o_{2}}^{s_{1},s_{2}}=c^{\dagger}_{o_{2},s_{2}}(\tau,\boldsymbol{r})c_{o_{1},s_{1}}(\tau,\boldsymbol{r})$ and
$n_{o_{1}}^{s_{1}}=n_{o_{1},o_{1}}^{s_{1},s_{1}}=c^{\dagger}_{o_{1},s_{1}}(\tau,\boldsymbol{r})c_{o_{1},s_{1}}(\tau,\boldsymbol{r})$. In principle, the two approaches should yield compatible results as long as the underlying assumptions of the approach are valid. Furthermore, it should be noted that if the initial model contains no inter-orbial hopping between orbitals which are interacting, the renormalizations of $\tilde{P}$ for both TPSC3 and TPSC5 as well as $\tilde{C}^{\uparrow,\downarrow}$ in TPSC5 are singular and thus the local and static limits cannot be captured in the standard fashion. Therefore, the question arises on how to renormalize these components in such a case. The first option is to stay at the level of plain Hartree-Fock which does not renormalize these couplings leading to an overestimation of their contribution. The second option is to use the freedom of the Ansatz and employ the same rescaling as we use for the other components. In principle, as long as the trace consistency check (equivalence of left and right hand side in Eq. (10)) is fulfilled, both variants are reasonable. However the former is expected to break down rapidly at stronger coupling, while the latter maintains the Kanamori-Brückner-like scaling [80]. We will maintain the first variant for easier comparability with earlier multi-orbital TPSC approaches [89, 86].
@@ 3 Benchmarking
$H^{HK}=U\hat{n}_{o_{1^{\prime}}}^{\uparrow}\hat{n}_{o_{1^{\prime}}}^{\downarrow}+\sum_{{o_{1^{\prime}}}\neq{o_{2^{\prime}}}}$ $\left((U-2J)\hat{n}_{o_{1^{\prime}}}^{\uparrow}\hat{n}_{o_{2^{\prime}}}^{\downarrow}+(U-3J)\hat{n}_{o_{1^{\prime}}}^{s^{\prime}}\hat{n}_{o_{2^{\prime}}}^{s^{\prime}}\right.$ (47)
$\quad\left.-J\hat{c}^{{\dagger}}_{o_{1^{\prime}},\uparrow}\hat{c}_{o_{1^{\prime}},\downarrow}\hat{c}^{{\dagger}}_{o_{2^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\uparrow}+J\hat{c}^{{\dagger}}_{o_{1^{\prime}},\uparrow}\hat{c}^{{\dagger}}_{o_{1^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\downarrow}\hat{c}_{o_{2^{\prime}},\uparrow}\right)\,.$
@@ 3.2 Analysis of the model with $J=U/3$
From these observations, we can extract a guiding principle of when to expect TPSC to work reliably in a multi-orbital setting and when not: If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than if the states are close in energy. Furthermore, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. This is intrinsically tied to Hartree-Fock being essentially a ground state targeting approach which performs better at lower temperatures. However, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC typically fails at low temperatures.
@@ Appendix C Derivation of the interaction from functional derivatives
Here the first term in the second line vanishes due to $SU(2)$ symmetry. Note that if we would have spin-orbit coupling the time-reversal symmetry would not lead to a cancellation of this contribution, in contrast to the single-band case [107, 87]. The second contribution reads
@@ Appendix E Comparison to ED
Figure 6: Comparison between TPSC3 (blue), TPSC5 (yellow) and ED (red) for the model defined in Eq. (144) at $\beta=0.5$ on the left and $\beta=2$ on the right, for different $U$ and $J$ combinations. We include the zeroth order self-energy in both TPSC variants.

### v5

@@ 2 Multi-orbital TPSC
$=\Big{(}D_{o_{4^{\prime}},o_{1}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$ (8)
$\quad+C_{o_{1},o_{3^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}$
$\quad+P_{o_{2^{\prime}},o_{1}}\braket{T_{\tau}c^{\dagger}_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}\Big{)}\,.$
@@
$\Big{(}D_{o_{4^{\prime}},o_{1}}\braket{T_{\tau}c_{o_{4^{\prime}},s_{4^{\prime}}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})}$
$\quad-C_{o_{1},o_{3^{\prime}}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3^{\prime}}s_{1}}(\boldsymbol{\tau}_{1})c_{o_{1},s_{1}}^{\dagger}(\boldsymbol{\tau}_{1})c_{o_{1},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})}$
@@
By combining Eq. (9) and Eq. (10) we obtain the explicit form of the Ansätze. It should be noted that there is an ambiguity in the way we define the prefactors: we can either define them as objects depending on one or on two spin indices. In the former case, we end up with three independent vertices, while in the latter case, we end up with five. (The ansatz proportional to $P$ always depends only on a single spin.) The origin of this ambiguity is the freedom of pulling the sum over the internal spin in Eq. (8) into the Ansatz (leading to TPSC3) or performing the Ansatz inside the sum (leading to TPSC5). Whether both of these choices are valid in the sense that no artificial $SU(2)$ symmetry breaking is introduced has to be checked by deriving the interaction vertex induced by them. Again we stress that there is technically a much larger group of possible Ansätze due to the freedom of shifting contributions to different interaction terms.
@@
In these equations, we introduced a short hand notation which allows us to drop redundant arguments, i.e. $n_{o_{1},o_{2}}^{s_{1},s_{2}}=c^{\dagger}_{o_{2},s_{2}}(\boldsymbol{\tau})c_{o_{1},s_{1}}(\boldsymbol{\tau})$ and $n_{o_{1}}^{s_{1}}=n_{o_{1},o_{1}}^{s_{1},s_{1}}=c^{\dagger}_{o_{1},s_{1}}(\boldsymbol{\tau})c_{o_{1},s_{1}}(\boldsymbol{\tau})$. Note that we dropped the time and position argument as $n$ only depends, due to translational invariance, on the difference between the arguments on the right hand side, which are identical. In principle, the two approaches should yield compatible results as long as the underlying assumptions of the approach are valid. Furthermore, it should be noted that if the initial model contains no inter-orbial hopping between orbitals which are interacting, the renormalizations of $\tilde{P}$ for both TPSC3 and TPSC5 as well as $\tilde{C}^{\uparrow,\downarrow}$ in TPSC5 are singular and thus the local and static limits cannot be captured in the standard fashion. Therefore, the question arises on how to renormalize these components in such a case. The first option is to stay at the level of plain Hartree-Fock which does not renormalize these couplings leading to an overestimation of their contribution. The second option is to use the freedom of the Ansatz and employ the same rescaling as we use for the other components. In principle, as long as the trace consistency check (equivalence of left and right hand side in Eq. (10)) is fulfilled, both variants are reasonable. However the former is expected to break down rapidly at stronger coupling, while the latter maintains the Kanamori-Brückner-like scaling [80]. We will maintain the first variant for easier comparability with earlier multi-orbital TPSC approaches [89, 86].
@@ 3 Benchmarking
$H^{HK}=U\hat{n}_{o_{1^{\prime}}}^{\uparrow}\hat{n}_{o_{1^{\prime}}}^{\downarrow}+\sum_{{o_{1}}\neq{o_{2}}}$ $\left((U-2J)\hat{n}_{o_{1}}^{\uparrow}\hat{n}_{o_{2}}^{\downarrow}+(U-3J)\hat{n}_{o_{1}}^{s^{\prime}}\hat{n}_{o_{2}}^{s^{\prime}}\right.$ (47)
$\quad\left.-J\hat{c}^{{\dagger}}_{o_{1},\uparrow}\hat{c}_{o_{1},\downarrow}\hat{c}^{{\dagger}}_{o_{2^{\prime}},\downarrow}\hat{c}_{o_{2},\uparrow}+J\hat{c}^{{\dagger}}_{o_{1},\uparrow}\hat{c}^{{\dagger}}_{o_{1},\downarrow}\hat{c}_{o_{2},\downarrow}\hat{c}_{o_{2},\uparrow}\right)\,.$
@@ 3.2 Analysis of the model with $J=U/3$
From these observations, we can extract a guiding principle of when to expect TPSC to work reliably in a multi-orbital setting and when not: If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than if the states are close in energy. Furthermore, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. This is intrinsically tied to Hartree-Fock being essentially a ground state targeting approach which performs better at lower temperatures. However, at very low temperatures, we expect the envelope diagram to give a sizable contribution near the Fermi-level inducing a non-local and non-static irreducible Vertex [107, 108]. Additionally, it was shown in an RG analysis that in the low temperature limit, the vertex is only RG relevant at the Fermi-surface again implying a momentum and Frequency dependence strongly confined in both spaces [28]. Therefore, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC is expected to fail at very low temperatures.
@@ 3.3.1 Comparison to DMFT - TPEV’s
Figure 5: Comparison of $\Sigma_{1,1}$ at different momentum points (indicated in the bottom right corner) as a function of Matsubara frequency between TPSC3 (blue), TPSC5 (orange), TPSC from Ref. [90] (green) and D-TRILEX (purple) at $U=2$ and $J=0.25U$. The inverse temperature is fixed to $\beta=2$.
@@ 3.3.2 Comparison to D-TRILEX - self-energy and susceptibilities
An analogous picture emerges when comparing the self-energy calculated with the different methods at different momentum points as a function of Matsubara frequency, see Fig. 5. We observe qualitative agreement but quantitative deviations. Note that at high frequency the noise of the DMFT impurity solver utilized to obtain the starting point in D-TRILEX becomes observable, especially in the real part of the self-energy.
@@ Appendix C Derivation of the interaction from functional derivatives
Here the first term in the second line vanishes due to $SU(2)$ symmetry. Note that if we would have spin-orbit coupling the time-reversal symmetry would not lead to a cancellation of this contribution, in contrast to the single-band case [87]. The second contribution reads
@@ Appendix E Comparison to ED
Figure 8: Comparison between TPSC3 (blue), TPSC5 (yellow) and ED (red) for the model defined in Eq. (144) at $\beta=0.5$ on the top and $\beta=2$ on the bottom, for different $U$ and $J$ combinations. We include the zeroth order self-energy in both TPSC variants.